OpenStudy (anonymous):

Use the discriminant to find the number of real solutions of the equation 3x2 – 5x + 4 = 0. Show your work.

OpenStudy (anonymous):

3^2-5x+4=0

OpenStudy (anonymous):

I mean 3x^2

OpenStudy (abmon98):

a quadratic equation is in the form of ax^2+bx+c=0. The discriminant is determined from the coeffcients of your quadratic equation. D=b^2-4ac

OpenStudy (abmon98):

a=3 b=-5 c=4

OpenStudy (abmon98):

b^2-4ac < 0 means there are no real roots b^2-4ac = 0 means equal real roots b^2-4ac > 0 means there are distinct real roots

OpenStudy (anonymous):

wat to do next

OpenStudy (abmon98):

use the quadratic formula to find the roots of your quadratic equation but before doing that calculate your discriminant.

OpenStudy (abmon98):

b^2-4ac (-5)^2-4(3)(4)=-23 -b+-(b^2-4ac)^1/2/2a -5+(-23)^1/2/2*3 or -5-(23)^1/2/2(3)

OpenStudy (anonymous):

is that right Use the discriminant to find the number of real solutions of the equation 3x2 – 5x + 4 = 0. Show your work. b^2-4ac (-5)^2-4(3)(4)=-23 -b+-(b^2-4ac)^1/2/2a -5+(-23)^1/2/2*3 or -5-(23)^1/2/2(3)